A 0.145 kg baseball pitched at 39.0 m/s is hit on a horizontal line drive straight back toward the pitcher at 52.0 m/s. If the contact time between the bat and ball is 1.0 x 10^-3 s, calculate the average force between the ball and bat during contact.


1
Expert's answer
2021-10-18T11:10:22-0400

The average force can be found in terms of impulse:


"F=\Delta p/t=(mv_f-mv_i)/t,\\
F=m(v_f-v_i)/t,\\
F=0.145(52-(-39))/0.001=13195\text{ N}."